# Machine Learning for Synthetic Data Generation: A Review

### v8

@@ V-B Privacy Protection in Data Synthesis
Solutions have been proposed in two broad categories. In the first category, different data anonymization-based approaches such as K-anonymity [180, 181, 182] and nearest marginal [183] to sanitize data so that it cannot be easily re-identified. However, these data anonymization approaches do not provide rigorous privacy guarantees [14]. In the second category, synthetic data generation approaches have been proposed to generate realistic synthetic data using rigorous differential privacy definitions [165, 113, 184] for various applications. In particular, Bindschaedler et al. [165] introduced the idea of plausible deniability instead of directly adding noise to the generative model. This mechanism results in input indistinguishability that means by observing the output set (i.e., synthetics) an adversary cannot make sure whether a particular data record was in the input set (i.e., real data). With the help of generative modeling, Acs et al. [43] clusters the original datasets into k clusters with differentially private kernel k-means and produce synthetic data for each cluster. By comparison, Liu et al [166] introduce two-level privacy-preserving synthetic data generation. At the data level, a selection module is used to select the items which contribute less to the user’s preference. At the item level, a synthetic item generation module is developed to create the corresponding synthetic item.
@@
Taking advantage of the GAN, several methods are proposed to generate synthetic data to get better effect [167, 168, 49, 169] which closely matches the distribution of the source data than the hidden Markov model-based approach [71], RBF based approach [170] and Auto-encoder based approach [14]. Xie et al [169] propose DPGAN by adding noise on the gradient of the Wasserstein distance with respect to the training data. This approach does not adopt the optimization strategy to improve the training stability and convergence speed. To address these problems, Zhang et al. [167] proposed dp-GAN, a general private data publishing framework for rich semantic data without the requirement of tag information compared to [168]. By comparison, Beaulieu-Jones et al. [168] trained the discriminator under differentially private SGD, which generates plausible individuals of clinical datasets. Tseng and Wu [171] apply compressive privacy [185] for CPGAN, which would generate compressing representations that retain high utility. Jordon et al.  [49] modifies the Private Aggregation of Teacher Ensembles (PATE) framework and applies it to GANs. The proposed approach allows a tight bound on the influence of any individual sample on the model, resulting in tight differential privacy guarantees and thus an improved performance over models for data synthesis. While most of these approaches inject noise into the energy function, a differentially private GAN called GANobfuscator [172] achieve differential privacy by adding noise within the training procedure.
@@
While centralized differential privacy assumes data aggregators are reliable, local differential privacy (LDP) [184] assumes that aggregators cannot be trusted and relies on data providers to perturb their own data and is used to generate private synthetic datasets that is similar to the private dataset. [173] is inspired by PriView [174] but for computing any k-way marginals under the LDP setting for the marginal table release problem. Apart from LDP in distributed setting, Triastcyn and Faltings [175] propose federated generative privacy that utilizes insuffient local data from multiple clients to train a GAN. The method shares only generators that do not come directly into contact with data and the discriminator remain private. This model can output artificial data, not belonging to any real user in particular, but coming from the common cross-user data distribution.
@@
These privacy-preserving data synthesis methods mainly aim at structured data like tables, which cannot be applied to high dimensionality and complexity. To solve this problem, PriView [174] constructs the private k-way marginal tables for $k\geq 3$ by first extracting low-dimensional marginal views from the flat data and adding noise to the views and then applying a reprocessing technique to ensure the consistency of the noisy views. Zhang et al. [176] consider repetitive perturbation of the original data as a substitute to the original data with a synthetic data generation technique called PrivBayes. PrivBayes decomposes high dimensional data into low dimensional marginals by constructing a Bayesian network and injects noise into these learned low dimensional marginals to ensure differential privacy and the synthetic data is inferred from these noised marginals. Instead of the Bayesian network, differentially private auto-encoder [14] significantly improves the effectiveness of differentially private synthetic data release. [177] applies data cleaning method [186] to fix the violations on the structure of the data in the synthetic data. Instead of using graphical models as the summarization/representation of a dataset [14, 165, 187, 188, 167][178] proposes to use a set of large number of low-degree marginals to represent a dataset. The advantage of this approach is that it makes weak assumptions about the conditional independence among attributes, and simply tries to capture correlation relationships that are in the dataset. Meanwhile, the method is especially attractive under differential privacy for its straightforward sensitivity measurement, reduced noise variance, and efficient privacy cost.
@@ V-C Privacy Threats in Foundation Models
Entering the era of foundation models, recent research has demonstrated that training data can be exposed from large language models [189] as well as stable diffusion [190]. In both types of models, attackers can generate sequences from the trained model and identify those memorized from the training set. Studies have shown that a sequence that appears multiple times in the training data is more likely to be generated than a sequence that occurred only once [191, 192, 154]. Accordingly, Kandpal et al [193] propose to deduplicate the training data that appears multiple times such that the privacy risks in language models is mitigated.
@@
V-D Other strategies
Figure 3: Synthetic data applications
Other strategies include
1. Data Anonymization: This process removes personally identifiable information from data sets, ensuring that the individuals the data describe remain anonymous. This is crucial in industries such as healthcare, where patient data privacy is a legal requirement.
2. Data Masking: This technique involves replacing sensitive data with fictitious yet realistic data. It is often used to protect the data while maintaining its usability for testing or development purposes. For example, a developer might need to use customer data to test a new feature, but they don’t need to know the actual personal details of the customer to do so.
3. Data Perturbation: This involves adding noise to the data to prevent the identification of individuals in the dataset while preserving the statistical properties of the data. This is particularly useful in research scenarios where data needs to be shared but individual privacy must be maintained.
4. Differential Privacy: This system publicly shares information about a dataset by describing the patterns of groups within the dataset while withholding information about individuals in the dataset. It provides a mathematical guarantee of privacy and is becoming an increasingly popular method for enhancing privacy in machine learning.
5. Federated Learning: This is a machine learning approach where the model is trained across multiple decentralized edge devices or servers holding local data samples, without exchanging them. This approach allows for the utilization of a wide array of data sources, while also ensuring that sensitive data does not leave its original device.
These techniques are all part of the broader field of synthetic data generation, which aims to create data that can be used for a variety of purposes (such as training machine learning models) without compromising privacy.
@@ VI Fairness
Generating synthetic data that reflect the important underlying statistical properties of the real-world data may also inherit the bias from data preprocessing, collection, and algorithms [194]. The fairness problem is currently addressed by three types of methods [195]: (i) preprocessing, which revises input data to remove information correlated to sensitive attributes, usually via techniques like massaging, reweighting, and sampling. (ii) in-processing, which adds fairness constraints to the model learning process; and (iii) post-processing, which adjusts model predictions after the model is trained.
Most existing fairness-aware data synthesis methods leverage preprocessing techniques. The use of balanced synthetic datasets created by GANs to augment classification training has demonstrated the benefits for reducing disparate impact due to minoritized subgroup imbalance [196, 197, 198][199] models bias using a probabilistic network exploiting structural equation modeling as the preprocessing to generate a fairness-aware synthetic dataset. Authors in [200] leverage GAN as the pre-processing for fair data generation that ensures the generated data is discrimination free while maintaining high data utility. By comparison, [201] is geared towards high dimensional image data and proposes a novel auxiliary classifier GAN that strives for demographic parity or equality of opportunity. However, preprocessing would require the synthesized data provider to know all correlations, biases, and distributions of variables in the existing datasets as a priori. Compared to preprocessing, the latter two categories are less-developed for fair data synthesis.
@@
In the emerging AIGC using foundation models, the generated images and texts may also inherit the stereotypes, exclusion and marginalization of certain groups and toxic and offensive information in the real-world data. This would lead to discrimination and harm to certain social groups. The misuse of such data synthesis approaches by misinformation and manipulation would lead to further negative social impact [204]. Given that the quality of the data generated by foundation models is inextricably linked to the quality of the training corpora, it is essential to regulate the real-world data being used to form the data synthesis distribution. While reducing bias in data is important, the remaining bias in the data may also be amplified by the models [195] or the privacy-enhancing components [202]. With frequent inspection and sensitive and toxic information removal on both data and model, it will help govern the information generated from those foundation models and ensure the models would do no harm (we hope!).
VII Trustworthiness
As data-driven decision making proliferates across various industries, the creation and use of synthetic data has become increasingly prevalent. Synthetic data, artificially generated data that simulates real-world scenarios, offers a way to bypass several problems associated with real data, such as privacy concerns, scarcity, or data collection difficulty. Nevertheless, the trustworthiness of synthetic data is a subject of ongoing debate, hinging on aspects such as data representativeness, privacy preservation, and potential biases.
For synthetic data to be trustworthy, it must offer a faithful statistical representation of the original data, while maintaining the inherent variability and structure. The risk lies in creating data that oversimplifies or misrepresents the complexities of real-world data, potentially leading to inaccurate conclusions or ineffective solutions when used in analysis or modelling.
Privacy preservation is another critical factor in synthetic data generation. Synthetic data is often utilized in situations where the use of real data may breach privacy regulations or ethical boundaries. While synthetic data promises a level of anonymity, there is an ongoing debate about the extent to which this data can be de-anonymized. If synthetic data could be traced back to the original contributors, it would undermine its trustworthiness and the privacy it promises to uphold.
Potential biases in synthetic data are a significant concern. Even though synthetic data is artificially generated, it often relies on real-world data to inform its creation. Thus, if the real-world data is biased, these biases could be unwittingly replicated in the synthetic data, perpetuating the same flawed patterns and undermining its trustworthiness.
Moreover, assessing the trustworthiness of synthetic data involves the evaluation of the synthetic data generation methods themselves. Transparency in the generation process, including a clear understanding of the underlying algorithms and parameters used, is crucial in judging the trustworthiness of the resultant synthetic data.
In conclusion, while synthetic data presents compelling benefits, the trustworthiness of such data depends on its representativeness, privacy preservation, and absence of bias. By recognizing and addressing these concerns, researchers and practitioners can make informed decisions about synthetic data’s validity and ethical implications. Transparent and robust synthetic data generation methods are paramount in fostering this trust.
VIII Evaluation Strategy

### v9

@@ V-B Privacy Protection in Data Synthesis
Solutions have been proposed in two broad categories. In the first category, different data anonymization-based approaches such as K-anonymity [217, 218, 219] and nearest marginal [220] to sanitize data so that it cannot be easily re-identified. These data anonymization approaches involve replacing sensitive data with fictitious yet realistic data. It is often used to protect the data while maintaining its usability for testing or development purposes. However, they often do not provide rigorous privacy guarantees [14]. In the second category, synthetic data generation approaches have been proposed to generate realistic synthetic data using rigorous differential privacy definitions [182, 112, 221] for various applications. These approaches involves adding noise to the data to prevent the identification of individuals in the dataset while preserving the statistical properties of the data. This is particularly useful in scenarios where data needs to be shared but individual privacy must be maintained. In particular, Bindschaedler et al. [182] introduced the idea of plausible deniability instead of directly adding noise to the generative model. This mechanism results in input indistinguishability that means by observing the output set (i.e., synthetics) an adversary cannot make sure whether a particular data record was in the input set (i.e., real data). With the help of generative modeling, Acs et al. [43] clusters the original datasets into k clusters with differentially private kernel k-means and produce synthetic data for each cluster. By comparison, Liu et al [213] introduce two-level privacy-preserving synthetic data generation. At the data level, a selection module is used to select the items which contribute less to the user’s preference. At the item level, a synthetic item generation module is developed to create the corresponding synthetic item.
@@
Taking advantage of the GAN, several methods are proposed to generate synthetic data to get better effect [184, 191, 49, 185, 27, 195, 196] which closely matches the distribution of the source data than the hidden Markov model-based approach [71], RBF based approach [202], Bayesian network-based [212], and Auto-encoder based approach [14]. Xie et al [185] propose DPGAN by adding noise on the gradient of the Wasserstein distance with respect to the training data. This approach does not adopt the optimization strategy to improve the training stability and convergence speed. To address these problems, Zhang et al. [184] proposed dp-GAN, a general private data publishing framework for rich semantic data without the requirement of tag information compared to [191]. By comparison, Beaulieu-Jones et al. [191] trained the discriminator under differentially private SGD, which generates plausible individuals of clinical datasets. Tseng and Wu [183] apply compressive privacy [222] for CPGAN, which would generate compressing representations that retain high utility. Jordon et al.  [49] modifies the Private Aggregation of Teacher Ensembles (PATE) framework and applies it to the discriminator of GANs. The proposed approach perceives the discriminator as a classifier and utilizes its output as knowledge such that the student learns from noisy labels that are obtained through privately aggregating the discriminator’ votes. This allows a tight bound on the influence of any individual sample on the model, resulting in tight differential privacy guarantees and thus an improved performance over models for data synthesis. By comparison, Long et at. [197] applies teacher-student-based differential privacy to the generator. While most of these approaches inject noise into the energy function, a differentially private GAN called GANobfuscator [186] achieve differential privacy by adding noise within the training procedure.
@@
While centralized differential privacy assumes data aggregators are reliable, local differential privacy (LDP) [221] assumes that aggregators cannot be trusted and relies on data providers to perturb their own data and is used to generate private synthetic datasets that is similar to the private dataset. [215] is inspired by PriView [216] but for computing any k-way marginals under the LDP setting for the marginal table release problem. Furthermore, [198] considers DP at label-level on GAN for synthetic spatial point generation. Apart from LDP in distributed setting, Triastcyn and Faltings [187] propose federated generative privacy that utilizes insufficient local data from multiple clients to train a GAN. The method shares only generators that do not come directly into contact with data and the discriminator remain private. This model can output artificial data, not belonging to any real user in particular, but coming from the common cross-user data distribution.
@@
These privacy-preserving data synthesis methods mainly aim at structured data like tables, which cannot be applied to high dimensionality and complexity. To solve this problem, PriView [216] constructs the private k-way marginal tables for $k\geq 3$ by first extracting low-dimensional marginal views from the flat data and adding noise to the views and then applying a reprocessing technique to ensure the consistency of the noisy views. [223, 224, 225, 226] leverage copula functions for multi-dimensional differentially private synthesization. Zhang et al. [206] consider repetitive perturbation of the original data as a substitute to the original data with a synthetic data generation technique called PrivBayes. PrivBayes decomposes high dimensional data into low dimensional marginals by constructing a Bayesian network and injects noise into these learned low dimensional marginals to ensure differential privacy and the synthetic data is inferred from these noised marginals. Instead of the Bayesian network, differentially private auto-encoder [14] significantly improves the effectiveness of differentially private synthetic data release. [207] applies data cleaning method [227] to fix the violations on the structure of the data in the synthetic data. Instead of using graphical models as the summarization/representation of a dataset [14, 182, 208, 209, 184][210] proposes to use a set of large number of low-degree marginals to represent a dataset. The advantage of this approach is that it makes weak assumptions about the conditional independence among attributes, and simply tries to capture correlation relationships that are in the dataset. Meanwhile, the method is especially attractive under differential privacy for its straightforward sensitivity measurement, reduced noise variance, and efficient privacy cost. [199] leverages the Hermite polynomial features to encapsulate a higher degree of information within a smaller order of feature. [201] constructs a graph that explore pairwise dependence between attributes and applies the junction tree algorithm to obtain the Markov random field (MRF), from which the noisy marginals are generated and the synthetic data are sampled.
While private synthetic data generation algorithms are agnostic to downstream tasks, it is important to meet the utility requirements for downstream use. [228] proposes post-processing via resampling from the synthetic data to filter out samples that do not meet the selected utility measures, thus improving the utility of synthetic data.
@@ V-C Privacy Threats in Foundation Models
Entering the era of foundation models, recent research has demonstrated that training data can be exposed from large language models [229] as well as stable diffusion [230]. In both types of models, attackers can generate sequences from the trained model and identify those memorized from the training set. Studies have shown that a sequence that appears multiple times in the training data is more likely to be generated than a sequence that occurred only once [231, 232, 166]. Accordingly, Kandpal et al [233] propose to deduplicate the training data that appears multiple times such that the privacy risks in language models is mitigated. [234] is the first work to enforce privacy using differentially private stochastic gradient descent (DP-SGD) in diffusion models. Several attempts has been made to reduces the noise in the gradient during DP-SGD training and improves the generative quality in diffusion models, via semantic-aware pretraining [235, 236], latent information [237], and retrieval-augmented generation [238]. In the meantime, differential privacy has been heavily invested in privacy protection of large language models [239].
@@ VI Fairness
Generating synthetic data that reflect the important underlying statistical properties of the real-world data may also inherit the bias from data preprocessing, collection, and algorithms [240]. Minority groups can often end up being under-represented in synthetic data [241, 242, 243]. The fairness problem is currently addressed by three types of methods [244]: (i) preprocessing, which revises input data to remove information correlated to sensitive attributes, usually via techniques like massaging, reweighting, and sampling. (ii) in-processing, which adds fairness constraints to the model learning process; and (iii) post-processing, which adjusts model predictions after the model is trained.
Most existing fairness-aware data synthesis methods leverage preprocessing techniques. The use of balanced synthetic datasets created by GANs to augment classification training has demonstrated the benefits for reducing disparate impact due to minoritized subgroup imbalance [245, 246, 247][248] models bias using a probabilistic network exploiting structural equation modeling as the preprocessing to generate a fairness-aware synthetic dataset. Authors in [249] leverage GAN as the pre-processing for fair data generation that ensures the generated data is discrimination free while maintaining high data utility. By comparison, [250] is geared towards high dimensional image data and proposes a novel auxiliary classifier GAN that strives for demographic parity or equality of opportunity. However, preprocessing would require the synthesized data provider to know all correlations, biases, and distributions of variables in the existing datasets as a priori. Compared to preprocessing, the latter two categories are less-developed for fair data synthesis. [251] insert a structural causal model in the input layers of the generator, allowing each variable to be reconstructed conditioned on its causal parents for inference time debiasing.
@@
In the emerging AIGC using foundation models, the generated images and texts may also inherit the stereotypes, exclusion and marginalization of certain groups and toxic and offensive information in the real-world data. This would lead to discrimination and harm to certain social groups. The misuse of such data synthesis approaches by misinformation and manipulation would lead to further negative social impact [253]. Given that the quality of the data generated by foundation models is inextricably linked to the quality of the training corpora, it is essential to regulate the real-world data being used to form the data synthesis distribution. While reducing bias in data is important, the remaining bias in the data may also be amplified by the models [244] or the privacy-enhancing components [252]. With frequent inspection and sensitive and toxic information removal on both data and model, it will help govern the information generated from those foundation models and ensure the models would do no harm.
VII Evaluation Strategy
@@ VIII Challenges and Opportunities
Last but not the least, the rise of foundation models in data synthesis presents both significant challenges and opportunities. On one hand, foundation models can be exploited by malicious actors to create sophisticate jailbreak attacks, deepfakes, discrimination, exclusion and toxicity problems, misinformation harms, sensitive information disclosure, and malicious use. These models can generate human-like text and realistic images or videos, making it difficult for traditional security measures to detect malicious content. Furthermore, the accessibility and rapid advancement of these technologies lower the barrier for cybercriminals, enabling more sophisticated and widespread attacks. The ability to generate vast amounts of realistic, yet fake, data can also overwhelm and deceive traditional detection systems, leading to an increase in false negatives and undetected breaches. On the other hand, foundation models offer promising opportunities to bolster cybersecurity defenses. AI-driven anomaly detection systems can leverage generative models to simulate various attack scenarios, improving their ability to recognize and mitigate real-world threats. In the meantime, the quest for transparency and interpretability in generative models promotes research into explainable AI. By proactively addressing these machine learning risks, synthetic data generation can evolve to deliver more ethical, secure, and transparent solutions, ultimately harnessing its full potential to benefit society while mitigating its associated risks.